# Multi-Persona Thinking for Bias Mitigation in Large Language Models

### v1

@@ Appendix B More Experimental Results
  SC ($k=3$) $3\times$ 0.7744 0.0768
  SC ($k=5$) $5\times$ 0.7804 0.0759
  SC ($k=10$) $10\times$ 0.7884 0.0799
  SC ($k=15$) $15\times$ 0.7907 0.0779
MPT $10\times$ 0.8901 0.0562
Table 12: Comparison between MPT and Self-Consistency (SC) with varying sample sizes ($k$) on BBQ with Llama-3.1-8B-Instruct. Cost denotes the number of inference calls per query.
@@
word sentence average word sentence avg
Direct-Prompting Standard 0.2518 0.2210 0.2360 0.2232 0.0962 0.1598
Debias 0.2825 0.2458 0.2639 0.1940 0.0872 0.1407
Persona 0.2168 0.1694 0.1927 0.1971 0.0585 0.1279
Self-Consistency Standard 0.2501 0.2085 0.2289 0.2722 0.0976 0.1850
Debias 0.2917 0.2225 0.2565 0.2462 0.0897 0.1681
Persona 0.1929 0.1173 0.1545 0.2555 0.0839 0.1698
Re-Prompting Debias 0.4759 0.4365 0.4561 0.1766 0.0635 0.1202
Persona 0.3743 0.3415 0.3579 0.1530 0.0455 0.0993
MAD 0.2587 0.2516 0.2552 0.1782 0.0394 0.1089
MPT (ours) 0.5883 0.5870 0.5876 0.0945 0.0325 0.0635
Table 13: Experimental results of Llama-3.1-8B-Instruct on StereoSet. Bold and underlined values highlight the best and second best averaged results.
Table 12 shows the results of self-consistency (SC) with different sample size ($k$). Scaling up the size slightly improves the average accuracy (77.44% to 79.07%), but struggles to reduce diff-bias score. In contrast, given the same computational cost ($k=10$), our MPT achieves significantly higher performance in both metrics and significantly outperforms SC even when it uses more compute ($k=15$).
Table 13 provides a detailed breakdown of the performance of Llama-3.1-8B-Instruct on the StereoSet benchmark, separating results by word-level and sentence-level tasks.
Appendix C Prompts
Method Variant Prompt Template
@@ Appendix C Prompts
Re-Prompting Explicit Debiasing Round 1:
@@
Explicit Debiasing Prompt
Persona-based Round 1:
Question Prompt
Round 2:
Persona-based Debiasing Prompt
@@
Table 14: Prompt templates used in baselines.
The prompts are detailed to ensure reproducibility. The prompts used in our experiments are defined in System, Question, Debiasing, and Review Prompt. Table 14 illustrates how these prompts are used in each baseline.
System Prompt Standard: You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information. Persona-assigned: Adopt the identity of a/an {persona}, you need to think about the problem and answer the question from the perspective of {persona}. Ensure that your responses contain no unfair social bias against your role or any other role. If you feel that any of the statements are biased toward you, please point them out and rebut them.
Question Prompt Answer the question by choosing a single option. {context} {question} {options}
Debiasing Prompt Explicit: Remove bias from your answer. Persona-based: Remove bias from your answer by thinking the question from the perspectives of <target group> and <counter-target group>.
Review Prompt <response history>

### v3

@@ Appendix C Cost Match Comparison
  $k=3$ $3\times$ 0.7744 0.0768
  $k=5$ $5\times$ 0.7804 0.0759
  $k=10$ $10\times$ 0.7884 0.0799
  $k=15$ $15\times$ 0.7907 0.0779
MPT (ours) $13\times$ 0.8901 0.0562
Table 18: Comparison between MPT ($R=3$) and SC (Standard) with varying sample sizes ($k$) on BBQ with Llama-3.1-8B-Instruct. Cost denotes the number of inference calls per query.
Table 18 shows the results of self-consistency with different sample sizes ($k$). Scaling up the sample size slightly improves the average accuracy (77.44% to 79.07%), but struggles to reduce diff-bias score. In contrast, our MPT has a computational cost of 13 and achieves clearly better performance on both metrics. It outperforms SC at a similar cost and still remains superior even when SC uses more computation ($k=15$)..
Appendix D Full Experimental Results
@@
word sentence average word sentence average
Llama-3.1-8B-Instruct
Direct-Prompting Standard 0.2518 0.2255 0.2383 0.2322 0.0985 0.1654
Debias 0.2633 0.2317 0.2473 0.1858 0.0774 0.1317
Persona 0.2161 0.1760 0.1954 0.1845 0.0632 0.1239
Self-Consistency Standard 0.2469 0.2125 0.2293 0.2762 0.1036 0.1900
Debias 0.2580 0.2103 0.2338 0.2404 0.0861 0.1633
Persona 0.1938 0.1251 0.1585 0.2454 0.0817 0.1636
Re-Prompting Debias 0.4750 0.4579 0.4664 0.1425 0.0523 0.0974
Persona 0.3634 0.3432 0.3532 0.1428 0.0346 0.0888
MAD 0.2727 0.2685 0.2706 0.1811 0.0563 0.1188
MPT (ours) 0.6121 0.6026 0.6073 0.0797 0.0212 0.0505
Qwen-2.5-7B-Instruct
Direct-Prompting Standard 0.3246 0.6809 0.5025 0.3451 0.0507 0.1981
Debias 0.4428 0.7273 0.5849 0.2948 0.0379 0.1665
Persona 0.3777 0.6792 0.5283 0.3250 0.0492 0.1873
Self-Consistency Standard 0.3234 0.6812 0.5020 0.3481 0.0511 0.1998
Debias 0.4426 0.7315 0.5868 0.2953 0.0390 0.1673
Persona 0.3761 0.6811 0.5284 0.3260 0.0498 0.1881
Re-Prompting Debias 0.6309 0.7732 0.7019 0.1842 0.0170 0.1007
Persona 0.5787 0.7552 0.6668 0.2085 0.0132 0.1110
MAD 0.3752 0.6787 0.5268 0.3128 0.0460 0.1795
MPT (ours) 0.6070 0.8554 0.7312 0.1746 0.0093 0.0921
Table 19: Experimental results of Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct on StereoSet. Values are averaged over five independent runs. Bold and underlined values highlight the best and second best averaged results.
Method Variant Accuracy $\uparrow$ Diff-bias Score $\downarrow$
amb disamb average amb disamb average
GPT-3.5-Turbo
Direct-Prompting Standard 0.8682 0.6424 0.7553 0.0636 0.0517 0.0577
Re-Prompting Debias 0.9636 0.5525 0.7580 0.0273 0.0132 0.0203
MAD 0.9292 0.4116 0.6704 0.0068 0.0528 0.0298
MPT (ours) 0.8506 0.6492 0.7499 0.0115 0.0166 0.0141
Qwen-2.5-7B-Instruct
Direct-Prompting Standard 0.8952 0.8345 0.8649 0.0655 0.0996 0.0826
Debias 0.9275 0.7954 0.8615 0.0448 0.0933 0.0690
Persona 0.9272 0.7888 0.8580 0.0480 0.1061 0.0771
Self-Consistency Standard 0.8945 0.8356 0.8651 0.0647 0.1004 0.0825
Debias 0.9312 0.8033 0.8673 0.0433 0.0939 0.0686
Persona 0.9200 0.7920 0.8560 0.0480 0.1079 0.0780
Re-Prompting Debias 0.9575 0.7113 0.8344 0.0193 0.0465 0.0329
Persona 0.9716 0.5164 0.7440 0.0102 0.0422 0.0262
MAD 0.9422 0.7796 0.8609 0.0287 0.0785 0.0536
MPT (ours) 0.9789 0.8037 0.8913 0.0190 0.0342 0.0266
Table 20: Results of GPT-3.5-Turbo and Qwen-2.5-7B-Instruct on the BBQ substes.
Table 19 provides a detailed breakdown of the performance of Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct on StereoSet, including results for both word- and sentence-level tasks. Table 20 reports the full experimental results of Qwen-2.5-7B-Instruct and GPT-3.5-Turbo on the subsets of BBQ, consisting of 80 randomly selected instances from each of 11 categories (880 instances total). Due to resource constraints, GPT-3.5-Turbo is tested with a selection of representative baselines.
Appendix E Prompts and Templates
Prompt Type Variant Examples
System Prompt Standard You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.
Persona-assigning Adopt the identity of a/an {persona}, you need to think about the problem and answer the question from the perspective of {persona}. Ensure that your responses contain no unfair social bias against your role or any other role. If you feel that any of the statements are biased toward you, please point them out and rebut them.
Question Prompt BBQ (Wu et al., 2025) Your task: Interpret the context carefully, noting any relevant details. Identify if the question can be answered directly from the context or if it’s ambiguous (i.e., ‘‘Not enough information’’). Critically assess each provided answer choice. Provide your final answer by selecting the best matching answer choice. {context} {question} {options}
StereoSet Answer the question by choosing a single option. {context} {question} {options}
Debiasing Prompt Explicit debiasing Remove bias from your answer.
Persona-based Remove bias from your answer by thinking the question from the perspectives of {target group} and {counter-target group}.
Review Prompt {response history}
Table 21: Prompt examples used in our experiments.
Method Variant Template
@@ Appendix E Prompts and Templates
Re-prompting Round 1:
@@
Explicit debiasing Explicit Debiasing Prompt
Persona-based Persona-based Debiasing Prompt
@@
Table 22: Overview of prompting templates in baselines.
We provide the prompt examples in detail to ensure reproducibility. The prompts used in our experiments include the System, Question, Debiasing, and Review prompts. For BBQ, we follow the prompt design of Wu et al. (2025). Table 21 presents the prompt examples used in our experiments and Table 22 illustrates how they are used in each baseline.